TV broadcast antennas are the tallest artificial structures on Earth. In 1987, a 79-kg physicist placed himself and 392 kg of equipment at the top of a 556-m-high antenna to perform gravity experiments. By how much was the antenna compressed, if we consider it to be equivalent to a steel cylinder 0.33 m in radius?


1
Expert's answer
2021-03-26T11:41:22-0400

Compression:


"\Delta L=\epsilon L."

Strain:


"\epsilon=\frac \sigma E."

Stress under force F=(m+M)g:


"\sigma=\frac FA=\frac{(m+M)g}{\pi R^2}."

So, combine this to get the compression:


"\Delta L=L\frac{(m+M)g}{\pi R^2E}=3.75·10^{-5}\text{ m}."

*Young's modulus of steel is "E=2·10^{11}\text{ Pa}."